# Quantum Feature Space of a Qubit Coupled to an Arbitrary Bath

### v3

@@ 3 Experimental Results
3.1 Efficiency of Quantum Feature Space Derivation
To find the best fit solution for Eq. 19, we used the Moore-Penrose inverse [23]. We implemented the Moore-Penrose inverse and methodologies of Sec. 2.3 using Pytorch [24] and parallelised calculations using across a GPU. The GPU for all experiments and simulations was an NVIDIA GeForce RTX 3060 with 12GB of memory.
To test and verify the methods of Sec. 2.3 we used ground truth expectations and the corresponding $V_{O}$ operators from QDataSet [25]. A batch of 1000 expectations was used to derive the parameters $\mu_{O}$, $\theta_{O}$, and $\psi_{O}$, with a total time of $0.52\pm 0.05$ ms.
@@ 3.2 Quantum Simulation Implementation
Our quantum simulator followed the formalism defined in Sec. 2 and was implemented in Pytorch [24]. We noted inefficiencies in the simulator of [25], which led to long execution times, motivating the need for a redesigned simulator to allow frequent and fast experiments. In the redesigned simulator, an overall improvement was using tensor operations whenever possible for the operations presented in Sec. 2.
Our simulator further reduced execution time by computing the unitary at timestep $t$ via left multiplying the unitary at time step $t-1$ by the exponentiated Hamiltonian at time step $t$. This method was not used in the previous simulator to compute unitaries, and instead, unitaries were computed via recomputing the entire product of required exponentiated Hamiltonians.
A further improvement was that our simulation did not compute the toggled system bath unitary ($U_{\mathrm{SB}}^{\mathrm{tog}}(t)$) for all timesteps; instead, only the final time step toggled system bath unitary was calculated as this was the only unitary required for expectation calculations. By enforcing this, we could exploit the associativity of matrix multiplication and implement a binary tree structure to compute the final unitary, which required $O(\log(N))$ multiplications, where $N$ is the number of time steps.
We benchmarked our simulator and found what required a few days for single qubit simulation in the previous implementation [25] took 10 minutes with the new simulator, giving over a 400$\times$ improvement. Greater speed-ups can be achieved if computing at lower precision and or using a GPU with more memory, where both options enable larger batch sizes.
Note, for simulator experiments, we derived the quantum feature space parameters directly and did not calculate expectations as we had access to the $Q_{O}D_{O}Q_{O}^{\dagger}$ operators. However, when using expectations from a physical quantum computer, one only needs to calculate $U_{\mathrm{ctrl}}(T)\rho U_{\mathrm{ctrl}}^{\dagger}(T)$, for the final time step $T$ (using the efficient binary tree structure), and then solve Eq. 19, using efficient methods such as the parallelised Moore-Penrose inverse, to find the parameters $\alpha_{O}$, $\beta_{O}$, and $\gamma_{O}$.
3.3 Creation of Quantum Feature Space Data
To create the quantum feature space data, we started with noise models and realisation generation methods following that of QDataSet [25]. All noise processes acted along the $x$ and $z$ axes, with the noise processes being correlated between the axes. This correlation was achieved by setting the noise values along the $z$-axis to be the absolute value of the $x$-axis values (similar to noise profile N6 in QDataSet).
@@ 3.3 Creation of Quantum Feature Space Data
There were six different noise processes in total. The first three were the above noise processes, and the other three were non-stationarity variants of each noise process. To achieve non-stationarity, we multiplied the noise processes by a deterministic triangular time envelope. The envelope was a triangular function with a peak at $T/2$ and a width of $T$.
The second noise process ‘$1/f$ + bump’ was chosen because it shared similarities with the first noise process ($1/f$). It was hypothesised that these two noise processes would result in parameters close together in the quantum feature space. Conversely, the third process (Gaussian coloured) was chosen because it was hypothesised that it would result in parameters far from the points of the first two noise processes.
@@
where $\sigma=\frac{T}{\lambda M}$, $\lambda\in\mathbb{R}^{+}$ is the pulse width, $T$ is the total time, $M$ is the number of time steps, $\tau_{n}=\left(\frac{n-0.5}{n_{\max}}\right)T+\delta_{\tau}$, with $\delta_{\tau}\in\mathbb{R}$ being used to add jitter to the pulse timing, and $A_{n}$ is the amplitude of the $n^{\mathrm{th}}$ pulse. For all experiments, the number of pulses was $n_{\max}=5$, 1024 discrete time steps were used and ensemble averaging was performed over 2000 realisations of the noise process for a given control pulse sequence.
3.4 Black Box Classification of Bath Through Quantum Feature Space
@@
We conducted a series of simulations to test the ability of the quantum feature space to classify distinct noise processes, where the quantum feature space ($Q_{O}$) is the space formed by using the three parameters $\alpha_{O}$, $\beta_{O}$, and $\gamma_{O}$ from Eq. 16. We first simulated qubit dynamics subject to realisations of the above noise processes with the approximated ideal CPMG pulses. We then simulated a ‘$1/f$ + bump’ process changing the location of its peak compared to earlier noise processes, which emulated an unknown noise process from a physical qubit that an experimentalist would be trying to classify.
@@
While simulating a qubit in the presence of the realisation of an unknown noise process, we used more realistic CPMG pulses; the pulse width was increased, and jitter was added to the peak’s timing and value. From the realistic CPMG pulses, we ended up with a cluster of points in the feature space arising from distinct realisations of the jitter. In a black box manner, we identify which noise process characterises the unknown noise process by measuring which a priori noise process the cluster of points was closest to in the quantum feature space.
@@
To simulate ideal CPMG pulses, we set $\lambda=\frac{1}{96}$, $\delta_{\tau}=0$, $n_{\max}=5$, and $A_{n}=\pi$. These values were chosen to approximate delta function pulses, shown in Fig. 2(a). For the realistic CPMG pulses, shown in Fig. 2(b), we set $\lambda=\frac{1}{24}$, $\delta_{\tau}$ was chosen from a uniform random distribution spanning the interval $[-\frac{24T}{M},\frac{24T}{M}]$, and $A_{n}=\pi+\epsilon$ where $\epsilon$ was drawn from a uniform random distribution spanning the interval $[-\frac{\pi}{5},\frac{\pi}{5}]$. This is s. Fifty such unique realistic control pulse sequences were generated with ensemble averaging over 2000 noise process realisations for each control pulse, resulting in a cluster of 50 unique points in the quantum feature space.
@@
Figure 2: Control pulses used in the simulations. (a) shows the ideal CPMG pulses and (b) shows the realistic CPMG pulses.
@@
From the first round of simulations, we calculated the average of the Euclidean distances between the realistic data points and the single data point from the a priori noise processes for the feature spaces associated with $Q_{X}$, $Q_{Y}$, and $Q_{Z}$, with the results shown in Table 1. We found that the ‘$1/f$ + bump’ noise process had the smallest total average Euclidean distance across all feature spaces, indicating the stationary ‘$1/f$ + bump’ noise process is likely closest to the unknown noise process.
@@
To refine our classification of the unknown noise process, using the results of the first iteration, we simulated several ‘$1/f$ + bump’ noise profiles, varying the location of the peak in the PSD. Specifically, peaks were positioned at predefined frequency bins, such as the 15th or 30th bin. The result shown in Table 2 indicated that the noise processes with peaks in the $240^{\mathrm{th}}$ and $120^{\mathrm{th}}$ bins were the closest matches, as measured by the average Euclidean distances in the quantum feature space to the unknown process, which peaks at the $200^{\mathrm{th}}$ bin.
@@
$V_{x}$ 0.398 0.368 0.379 0.381 0.828 0.924
$V_{y}$ 0.303 0.307 0.246 0.269 0.812 0.857
$V_{z}$ 0.290 0.297 0.304 0.293 0.687 0.621
Total 0.991 0.973 0.929 0.943 2.326 2.402
Table 1: Average Euclidean distances between realistic data points and the corresponding noise process values in the feature spaces of $Q_{X}$, $Q_{Y}$, and $Q_{Z}$. For each $V_{O}$ (row), the lowest (best) value is bolded. The total distance is the sum of the distances across all three feature spaces. The overall closest noise process is the ‘$1/f$ + bump’ noise process, with the smallest total distance as bolded. NS indicates non-stationary noise processes.
@@
Table 2: Average Euclidean distances between realistic data points and the corresponding noise process values in the feature spaces of $Q_{X}$, $Q_{Y}$, and $Q_{Z}$. For each $V_{O}$ (row), the lowest (best) value is bolded. The total distance is the sum of the distances across all three feature spaces. The overall closest noise process is the $240^{\mathrm{th}}$ noise profile, as indicated by the lowest total distance.
@@
It is interesting to note the significantly larger distances between the ‘$1/f$ + bump’ noise processes and the coloured noise processes. This enforces the hypothesis that the quantum feature space encodes the noise process characteristics, as such all the coloured noise processes are significantly further away from the ‘$1/f$ + bump’ noise processes than the ‘$1/f$ + bump’ noise processes are from each other. On the other hand, each noise process has a similar distance metric to its corresponding non-stationary noise process. This indicates that the non-stationary and stationary noise processes result in similar parameters in the quantum feature space, indicating that the quantum feature space is not as sensitive to non-stationarity.
@@
We also observed a relationship between the location of the simulated PSD peak and the average Euclidean distance in the quantum feature space to the unknown process (which has a peak at the $200^{\mathrm{th}}$ bin). Specifically, noise processes with peaks closer to the $200^{\mathrm{th}}$ bin, such as the $120^{\mathrm{th}}$ and $240^{\mathrm{th}}$ bin profiles, resulted in smaller distances in the feature space (0.767 and 0.666 respectively). Conversely, peaks further away, like the $15^{\mathrm{th}}$ or $480^{\mathrm{th}}$ bins, corresponded to significantly larger distances (0.903 and 1.422). This correlation demonstrates that the quantum feature space effectively encodes the key characteristics of the noise PSD and that the “distance” between noise processes in terms of their spectral properties (i.e. the location of the peak in the PSD) is meaningfully represented by distance in the quantum feature space. This finding further strengthens the hypothesis that the quantum feature space is a valuable tool for noise spectroscopy.
@@
The results in this section demonstrate that one could simulate a broad class of noise processes to derive quantum features and compare them to quantum features from experimental data with an unknown noise process. Once the general noise process is found, a binary search could be performed to estimate specific parameters of this unknown noise process by minimising distance in the quantum feature space.

### v4

@@ 3 Experimental Results
3.1 Efficiency of QFS Derivation
To find the best fit solution for Eq. 17, we used the Moore-Penrose inverse [penrose1955generalized]. We implemented the Moore-Penrose inverse and methodologies of Sec. 2.2 using Pytorch [paszke2019pytorch] and parallelised calculations using across a GPU. The GPU for all experiments and simulations was an NVIDIA GeForce RTX 3060 with 12GB of memory.
To test and verify the methods of Sec. 2.2 we used ground truth expectations and the corresponding $\tilde{O}$ operators from QDataSet [perrier2022qdataset]. A batch of 1000 expectations was used to derive the parameters $\alpha_{O}$, $\beta_{O}$, and $\gamma_{O}$, with a total time of $0.52\pm 0.05$ ms.
@@ 3.2 Quantum Simulation Implementation
Our quantum simulator followed the formalism defined in Sec. 2 and was implemented in Pytorch [paszke2019pytorch]. We noted inefficiencies in the simulator of [perrier2022qdataset], which led to long execution times, motivating the need for a redesigned simulator to allow frequent and fast experiments. In the redesigned simulator, overall improvement was achieved by using tensor operations whenever possible for the operations presented in Sec. 2.
Like [perrier2022qdataset], we used the Lie-Trotter product formula to estimate the time-evolution operator, that is,
$U(t)=\mathcal{T}e^{-i\int_{0}^{t}H(s)\,ds}\approx\prod_{j=0}^{N-1}e^{-iH(t_{j})\Delta t}$ (18)
where $N$ is the number of time steps, $t_{j}$ is the time at the $j^{\mathrm{th}}$ time step, and $\Delta T=\frac{T}{N}$ is the time step size.
Efficiency gains were made by implementing two major algorithmic changes. Firstly, for Eq. 6, the time-ordered control unitaries ($U_{\mathrm{ctrl}}(t)$) for all time steps needed to be calculated. To do this, first, we created an array of the exponentiated control Hamiltonians
$e^{-iH(0)\Delta t},e^{-iH(1)\Delta t},\ldots,e^{-iH(N-1)\Delta t}]$ (19)
Since the unitary at the $j^{\mathrm{th}}$ time step is the product of $j^{\mathrm{th}}$ and all previous exponentiated control Hamiltonians, our array of control unitaries,
$U_{\mathrm{ctrl}}(0),U_{\mathrm{ctrl}}(1),\ldots,U_{\mathrm{ctrl}}(N-1)]$ (20)
became
$e^{-iH(0)\Delta t},e^{-iH(1)\Delta t}e^{-iH(0)\Delta t},\ldots,e^{-iH(N-1)\Delta t}\ldots e^{-iH(1)\Delta t}e^{-iH(0)\Delta t}]$ (21)
Since matrix multiplication is associative, we computed the above array using a Prefix Scan [harris2007parallel] such that only $O(\log(N))$ iterations of matrix multiplications were required.
A further improvement was that our simulation did not compute the modified interaction unitary ($\tilde{U}_{\mathrm{I}}(t)$) for all timesteps; instead, only the final time step modified interaction unitary was calculated, as this was the only unitary required for expectation calculations. By enforcing this, again we could exploit the associativity of matrix multiplication and implement a binary tree structure to compute the final unitary, which also only required $O(\log(N))$ iterations of matrix multiplications.
[…]
Note, for simulator experiments, we derived the QFS parameters directly and did not calculate expectations as we had access to the $\tilde{O}$ operators. However, when using expectations from a physical quantum computer, one only needs to calculate $U_{\mathrm{ctrl}}(T)\rho U_{\mathrm{ctrl}}^{\dagger}(T)$, for the final time step $T$ (using the efficient binary tree structure), and then solve Eq. 17, using efficient methods such as the parallelised Moore-Penrose inverse, to find the parameters $\alpha_{O}$, $\beta_{O}$, and $\gamma_{O}$.
3.3 Creation of QFS Data
To create the QFS data, we started with noise models and realisation generation methods following that of QDataSet [perrier2022qdataset]. All noise processes acted along the $x$ and $z$ axes, with the noise processes being correlated between the axes. This correlation was achieved by setting the noise values along the $z$-axis to be the absolute value of the $x$-axis values (similar to noise profile N6 in QDataSet).
@@ 3.3 Creation of QFS Data
There were six different noise processes in total. The first three were the above noise processes, and the other three were non-stationary variants of each noise process. To achieve non-stationarity, we multiplied the noise processes by a deterministic triangular time envelope. The envelope was a triangular function with a peak at $T/2$ and a width of $T$.
The second noise process, ‘$1/f$ + bump’, was chosen because it shared similarities with the first noise process ($1/f$). It was hypothesised that these two noise processes would result in parameters close together in the QFS. Conversely, the third process (Gaussian coloured) was chosen because it was hypothesised that it would result in parameters far from the points of the first two noise processes.
@@
where $\sigma=\frac{T}{\lambda M}$, $\lambda\in\mathbb{R}^{+}$ is the pulse width, $T$ is the total time, $M$ is the number of time steps, $\tau_{n}=\left(\frac{n-0.5}{n_{\max}}\right)T+\delta_{\tau}$, with $\delta_{\tau}\in\mathbb{R}$ being used to add jitter to the pulse timing, and $A_{n}$ is the amplitude of the $n^{\mathrm{th}}$ pulse. For all experiments, the number of pulses was $n_{\max}=5$, 1024 discrete time steps were used, and ensemble averaging was performed over 2000 realisations of the noise process for a given control pulse sequence.
3.4 Black Box Classification of Bath Through QFS
@@
We conducted a series of simulations to test the ability of the QFS to classify distinct noise processes. We first simulated qubit dynamics subject to realisations of the above noise processes with the approximated ideal CPMG pulses. We then simulated a ‘$1/f$ + bump’ process, changing the location of its peak compared to earlier noise processes, which emulated an unknown noise process from a physical qubit that an experimentalist would be trying to classify (the unknown noise process had a peak at the $200^{\mathrm{th}}$ frequency bin). The goal was to classify the unknown noise process by comparing it to the known simulated noise processes in the QFS.
@@
While simulating a qubit in the presence of the realisation of an unknown noise process, we used more realistic CPMG pulses; the pulse width was increased, and jitter was added to the peak’s timing and value. From the realistic CPMG pulses, we ended up with a cluster of points in the feature space arising from distinct realisations of the jitter. In a black box manner, we identify which noise process characterises the unknown noise process by measuring which a priori noise process the cluster of points was closest to in the QFS.
@@
To simulate ideal CPMG pulses, we set $\lambda=\frac{1}{96}$, $\delta_{\tau}=0$, $n_{\max}=5$, and $A_{n}=\pi$. These values were chosen to approximate delta function pulses, shown in Fig. 2(a). For the realistic CPMG pulses, shown in Fig. 2(b), we set $\lambda=\frac{1}{24}$, $\delta_{\tau}$ was chosen from a uniform random distribution spanning the interval $\frac{24T}{M},\frac{24T}{M}]$, and $A_{n}=\pi+\epsilon$ where $\epsilon$ was drawn from a uniform random distribution spanning the interval $\frac{\pi}{5},\frac{\pi}{5}]$. This is s. Fifty such unique realistic control pulse sequences were generated with ensemble averaging over 2000 noise process realisations for each control pulse, resulting in a cluster of 50 unique points in the QFS.
@@
Figure 2: Control pulses used in the simulations. (a) shows the ideal CPMG pulses, and (b) shows the realistic CPMG pulses.
@@
From the first round of simulations, we calculated the average of the Euclidean distances between the realistic data points and the single data point from the a priori noise processes for the feature spaces associated with $\tilde{X}$, $\tilde{Y}$, and $\tilde{Z}$, with the results shown in Table 1. We found that the ‘$1/f$ + bump’ noise process had the smallest total average Euclidean distance across all feature spaces, indicating the stationary ‘$1/f$ + bump’ noise process is likely closest to the unknown noise process.
@@
To refine our classification of the unknown noise process, using the results of the first iteration, we simulated several ‘$1/f$ + bump’ noise profiles, varying the location of the peak in the PSD. Specifically, peaks were positioned at predefined frequency bins, such as the 15th or 30th bin. The result shown in Table 2 indicated that the noise processes with peaks in the $240^{\mathrm{th}}$ and $120^{\mathrm{th}}$ bins were the closest matches, as measured by the average Euclidean distances in the QFS to the unknown process, which peaks at the $200^{\mathrm{th}}$ bin.
To further refine our estimate of the unknown noise process, we conducted a third round of simulations, this time using the peaks between and including $130^{\mathrm{th}}$ and $230^{\mathrm{th}}$ bins. As the results show in Table 3, the noise processes with peaks at the $190^{\mathrm{th}}$ and $210^{\mathrm{th}}$ bins were the closest matches. If we then conclude that the unknown noise process is most similar to the ‘$1/f$ + bump’ processes with a peak at the $200^{\mathrm{th}}$ bin, we have then correctly estimated the unknown noise process.
@@
$\tilde{X}$ 0.40 0.37 0.38 0.38 0.83 0.92
$\tilde{Y}$ 0.30 0.31 0.25 0.27 0.81 0.86
$\tilde{Z}$ 0.29 0.30 0.30 0.29 0.69 0.62
Total 0.99 0.97 0.93 0.94 2.33 2.40
Table 1: Average Euclidean distances between realistic data points and the corresponding noise process values in the feature spaces of $\tilde{X}$, $\tilde{Y}$, and $\tilde{Z}$. For each $\tilde{O}$ (row), the lowest (best) value is bolded. The total distance is the sum of the distances across all three feature spaces. The overall closest noise process is the ‘$1/f$ + bump’ noise process, with the smallest total distance as bolded. NS indicates non-stationary noise processes.
@@
Table 2: Average Euclidean distances between realistic data points and the corresponding noise process values in the feature spaces of $\tilde{X}$, $\tilde{Y}$, and $\tilde{Z}$. For each $\tilde{O}$ (row), the lowest (best) value is bolded. The total distance is the sum of the distances across all three feature spaces. The overall closest noise process is the $240^{\mathrm{th}}$ noise profile, as indicated by the lowest total distance.
$130^{\mathrm{th}}$ $150^{\mathrm{th}}$ $170^{\mathrm{th}}$ $190^{\mathrm{th}}$ $210^{\mathrm{th}}$ $230^{\mathrm{th}}$
$\tilde{X}$ 0.318 0.301 0.305 0.289 0.273 0.267
$\tilde{Y}$ 0.155 0.127 0.101 0.081 0.074 0.111
$\tilde{Z}$ 0.297 0.282 0.303 0.290 0.278 0.282
Total 0.770 0.710 0.709 0.660 0.625 0.661
Table 3: Average Euclidean distances between realistic data points and the corresponding noise process values in the feature spaces of $\tilde{X}$, $\tilde{Y}$, and $\tilde{Z}$. For each $\tilde{O}$ (row), the lowest (best) value is bolded. The total distance is the sum of the distances across all three feature spaces. The overall closest noise process is the 210th noise profile, as indicated by the lowest total distance.
@@
It is interesting to note the significantly larger distances between the ‘$1/f$ + bump’ noise processes and the coloured noise processes. This enforces the hypothesis that the QFS encodes the noise process characteristics, as such, all the coloured noise processes are significantly further away from the ‘$1/f$ + bump’ noise processes than the ‘$1/f$ + bump’ noise processes are from each other. On the other hand, each noise process has a similar distance metric to its corresponding non-stationary noise process. This indicates that the non-stationary and stationary noise processes result in similar parameters in the QFS, indicating that the QFS is not as sensitive to non-stationarity.
@@
We also observed a relationship between the location of the simulated PSD peak and the average Euclidean distance in the QFS to the unknown process (which has a peak at the $200^{\mathrm{th}}$ bin). Specifically, noise processes with peaks closer to the $200^{\mathrm{th}}$ bin, such as the $190^{\mathrm{th}}$ and $210^{\mathrm{th}}$ bin profiles, resulted in smaller distances in the feature space (0.60 and 0.661, respectively). Conversely, peaks further away, like the $15^{\mathrm{th}}$ or $480^{\mathrm{th}}$ bins, corresponded to significantly larger distances (0.92 and 1.43). This correlation demonstrates that the QFS effectively encodes the key characteristics of the noise PSD and that the “distance” between noise processes in terms of their spectral properties (i.e. the location of the peak in the PSD) is meaningfully represented by distance in the QFS. This finding further strengthens the hypothesis that the QFS is a valuable tool for noise spectroscopy.
@@
The results in this section demonstrate that one could simulate a broad class of noise processes to derive quantum features and compare them to quantum features from experimental data with an unknown noise process. Once the general noise process is found, a binary search could be performed to estimate specific parameters of this unknown noise process by minimising distance in the QFS.
3.5 Explicit Comparison to Previous Work
The primary innovation of the QFS method is its dramatic reduction in model complexity and data requirements, directly addressing the key limitations of deep learning strategies [wise2021using, martina2023deep, youssry2020characterization]. As shown in Table 4, deep learning models employ neural networks with parameter counts in the tens or hundreds of thousands ($10^{4}$-$10^{5}$). Training these models demands a large number of examples (2100 - 3625 in the case of [youssry2020characterization]). For noise identification, the QFS method consists of just 10 parameters and in our proof-of-concept, we correctly identified an unknown noise process using only 18 simulated examples, representing a reduction of several orders of magnitude in both model complexity and data needs, making it a far more practical solution for real-time, adaptive characterisation.
In terms of experimental overhead, our method strikes a balance between simplicity and robustness. It requires more control sequences than the minimal approach of Fourier-Transform Noise Spectroscopy (FTNS) [vezvaee2024fourier], which can use a single free-induction decay. However, this trade-off yields significant resilience to experimental non-idealities. Our results were achieved using realistic CPMG pulses with inherent jitter and amplitude errors, demonstrating that the QFS is robust to the types of imperfections common in hardware. This practical robustness is a key advantage over methods that assume ideal conditions. Furhter, the FTNS method assumes stationary, Gaussian noise, which can be a significant limitation in practical applications. In contrast, our QFS method is designed to handle a broader class of noise processes, including non-stationary and coloured noise.
Furthermore, the QFS framework avoids the “black box” problem of deep learning models. Our results show a clear, interpretable link between distance in the QFS and the physical properties of the noise spectrum (i.e., the location of the PSD peak). This is a step beyond a simple classification label.
It is important to note that the goal of our QFS method differs from that of some prior work. Techniques like FTNS aim for a full, high-resolution reconstruction of a specific PSD. Our method, by contrast, is designed for efficient and robust classification and characterisation of the dominant noise family. As our results show, it can effectively identify the correct noise model and refine key parameters via a simple iterative search. While we observed that the current feature space is less sensitive to non-stationarity, its overall performance demonstrates that it fills a critical gap: providing a scalable, data-efficient, and computationally trivial tool for practical quantum noise analysis.
Attribute FTNS [vezvaee2024fourier] NN Regression [wise2021using, martina2023deep] Grey-box ML [youssry2020characterization] QFS (This Work)
Principle Analytical inversion (Fourier transform) Deep NN Deep NN + analytical layers Euclidean distance
Input Data Free-induction decay or single-pulse echo Coherence curves Observables expectations Observable expectations
Exp. Overhead One pulse One to several sequences Several thousand examples Tens of examples
Comput. Cost Analytical - very low NN training - high NN training - high Linear regression + distance calc - low
Key Advantage High spectral resolution High accuracy and robust denoising High accuracy on non-Markovian noise Data and computationally efficient, noise agnostic
Key Limitation Assumes stationary, Gaussian noise; sensitive to pulse errors Large datasets and ”black box” nature Large datasets required Not explicitly designed for PSD reconstruction
Model Parameters N/A $\sim 10^{4}-10^{5}$ (estimate) $\gtrsim 10^{5}$ (estimate) 10
Training Examples N/A Thousands Thousands 18
Table 4: Comparison of qubit noise spectrum estimation and classification methods. Our QFS approach is highlighted for its exceptionally low data and model complexity.